OpenStudy (anonymous):

Find the vertex of the graph of the function. f(x) = 2x2 - 8x + 9 ( 0, -1) ( -1, 0) ( 2, 1) ( 1, 2) I don't end up with either of the answers, can you explain how you find the answer but don't give me the answer

OpenStudy (whpalmer4):

You have a parabola:\[y =f(x)= ax^2+bx+c\]where \(a=2,\,b=-8,\,c=9\) The vertex of a parabola in that form has an x coordinate of \(-\dfrac{b}{2a}\) The y coordinate can be found by plugging in that value as \(x\) The correct answer does appear in the list :-)

OpenStudy (anonymous):

will -8 still be \[\frac{ -b }{ 2a }\]

OpenStudy (anonymous):

i meant will it still be a negative 8

OpenStudy (whpalmer4):

\[-\frac{b}{2a} = -\frac{-8}{2(2)} = \frac{8}{4}=\]

OpenStudy (anonymous):

That where i will get the 2 from but where would i get the second answer

OpenStudy (whpalmer4):

That's the x-coordinate of the vertex. You find the y-coordinate by evaluating \(y=f(x)\)